OpenStudy (anonymous):

What is an equation of the line in slope-intercept form? m = 2/7 and the y-intercept is (0, –12)

OpenStudy (anonymous):



OpenStudy (texaschic101):

y = mx + b slope (m) = 2/7 y intercept (b) = -12 just sub them in

OpenStudy (anonymous):

y=2/7x-12

OpenStudy (texaschic101):

correct